OpenStudy (anonymous):

\[(2-3\sqrt{-6})(-4+\sqrt{-6})\]

OpenStudy (anonymous):

same idea as before first write as \[(2-3\sqrt{6}i)(-4+\sqrt{6}i)\] and then multiply

OpenStudy (anonymous):

we can even do it the stupid "foil" way if you like it will work out the same

OpenStudy (anonymous):

hang on i got the answer already but i have another question

OpenStudy (saifoo.khan):

my calc's sayin MATH ERROR!

OpenStudy (anonymous):

ok yeah because either your calculator is lousy or you don't know how to enter a complex number

OpenStudy (anonymous):

\[(-4+\sqrt{-5})^{2}\]

OpenStudy (anonymous):

if you try to type in \[\sqrt{-6}\] it will tell you to go fish

OpenStudy (anonymous):

\[(-4+\sqrt{5}i)^2=(-4+\sqrt{5}i)(-4+\sqrt{5}i)\]

OpenStudy (anonymous):

lol, you guys are funny

OpenStudy (anonymous):

we get \[(-4\times -4-(\sqrt{5}\times \sqrt{5})+(-4\sqrt{5}-4\sqrt{5})i\]

OpenStudy (saifoo.khan):

Haha, yea.. It is saying me the same thing.

OpenStudy (anonymous):

better known as \[(16-5)+(-8\sqrt{5})i\] or \[11-8\sqrt{5}i\]

OpenStudy (anonymous):

steps clear?

OpenStudy (anonymous):

you're such a lifesaver! you're preventing me from taking another year of algebra II right now

OpenStudy (anonymous):

invoice is in the mail

OpenStudy (anonymous):

conjugate of 8+4i ?